A jeep is traveling at 50 km per hour when the brakes are applied. It comes to a halt after 11 seconds. Find the a) acceleration and b) the distance it has traveled after the brakes are applied.

 

 



1
Expert's answer
2021-09-20T09:59:50-0400

Given:

"v_0=50\:\rm km/h=13.9\: m/s"

"v=0\:\rm m/s"

"t=11\:\rm s"


(a) the acceleration

"a=\frac{v-v_0}{t}=\frac{0-13.9}{11}=-1.26\:\rm m/s^2"

(b) the distance it has traveled after the brakes are applied

"d=\frac{v+v_0}{2}t=\frac{0+13.9}{2}*11=76.5\:\rm m"